OpenStudy (anonymous):

Whats it the answer to 7 - 5/6 ?

OpenStudy (anonymous):

Use your calculator.

OpenStudy (anonymous):

wtf!

OpenStudy (anonymous):

\[\frac{7*6}{6} - \frac{5}{6} = \frac{37}{6}\]

OpenStudy (anonymous):

7 - 5/6 = ? hmmmm.....you must take the common denominator by treating the numbers as 7/1 - 5/6 = ? Common denominator is 6. Now, cross multiply and subtract.

OpenStudy (amistre64):

we now it needs to equal some number: 7 - 5/6 = N ; now clear the fraction 6(7 - 5/6 = N) 42 - 5 = 6N 37 = 6N ; and divide off that 6 N = 37/6